OpenStudy (anonymous):

What is the standard deviation of the data set given below? 5, 7, 9, 11, 13

OpenStudy (solomonzelman):

Did you find the mean (or the average) of your data?

OpenStudy (anonymous):

I think its 9

OpenStudy (sarahjoy26):

Hi @Gabe24Me ! Well, the group of numbers are 2 numbers apart. 5+2=7, 7+2=9, 9+2=11, ect.

OpenStudy (sarahjoy26):

Hop that helps a little bit.

OpenStudy (anonymous):

A. square root of 10 B. square root of 8 C. 3 D. square root of 3 E. 8 F. 10

OpenStudy (solomonzelman):

Yes, the mean is: \(\large\color{black}{ \displaystyle \mu=9 }\) Now, find \(\large\color{black}{ \displaystyle (x_i-\mu)^2 }\) That is \(\Downarrow\) Subtract the mean from each value in the list and square the difference

OpenStudy (sarahjoy26):

What is your explanation that the answer is a? @babylove18

OpenStudy (solomonzelman):

Then add all of those differences squared.

OpenStudy (anonymous):

ok thank you all for the help and A was correct

OpenStudy (solomonzelman):

I mean: \(\color{black}{ \displaystyle \frac{(5-9)^2+(7-9)^2+(9-9)^2+(11-9)^2+(13-9)^2}{5} }\) then, the population varience is: \(\color{black}{ \displaystyle \sigma_x{\tiny~}^2=\frac{(5-9)^2+(7-9)^2+(9-9)^2+(11-9)^2+(13-9)^2}{5} }\) then the population standard deviation: \(\color{black}{ \displaystyle \sigma_x=\sqrt{\frac{(5-9)^2+(7-9)^2+(9-9)^2+(11-9)^2+(13-9)^2}{5}} }\)

OpenStudy (solomonzelman):

sample standard deviation is same, but you divide by "1 less then the amount of data in the list", and in this case by 5-1=4. \(\color{black}{ \displaystyle s_x=\sqrt{\frac{(5-9)^2+(7-9)^2+(9-9)^2+(11-9)^2+(13-9)^2}{4}} }\)